OpenStudy (anonymous):

1. If 2x + 30 = 80, then x + 15 = 2. If 3x + 27 = 33, then x + 9 =

OpenStudy (freckles):

Solve 2x+30=80 for x

OpenStudy (freckles):

you could solve for x first then add 15 to it ...or

OpenStudy (freckles):

you could actually solve for x+15 if you wanted to since 2x+30=2(x+15) So how do you solve 2(x+15)=80 for x+15?